The main difference between aliphatic and aromatic hydrocarbons is that...

A. Aliphatic hydrocarbons can react to form products containing substituents such as Br and OH, whereas aromatic hydrocarbons cannot

B. Aromatic hydrocarbons have delocalized π electrons whereas aliphatic hydrocarbons do not

C. Aromatic hydrocarbons do not undergo any chemical reactions as they are chemically inert, but aliphatic hydrocarbons are chemically very reactive

D. Aromatic hydrocarbons only contain six carbon atoms, whereas aliphatic hydrocarbons can have more than six carbon atoms
Answer #1

The main difference between Aromatic hydrocarbon and Aliphatic hydrocarbon is

B. Aromatic hydrocarbons have delocalized π electrons whereas aliphatic hydrocarbons do not.

Aromatic Hydrocarbon has \pi electron ring system(e.g Benzene) which is delocalised around the enitire ring to give the resonance.A delocalized conjugated \pi system, most commonly an arrangement of alternating single and double bonds. It obeys Huckel's rule (4n+2)\pi rule.
Where as Aliphatic Hydrocarbon do not conjugated system and it carries linear chain , branched chain and do not have Conjugated system as well as do not obey the Huckel's rule (4n+2)\pi rule